# Solving Optimal Power Flow using a Variational Quantum Approach

### v1

@@ 6.2 Node permutation for qubit-efficient measurements
To solve the doubly parameterized OPF, we need to evaluate the Lagrangian function in (9) by measuring the observable in Lemma 1 so that we can compute its gradients over $(\boldsymbol{\theta},\alpha;\boldsymbol{\phi},\beta)$. We next adopt the XBM protocol to efficiently measure OPF observables.
@@
According to (17), observable matrix $\mathbf{M}$ has the $N\times N$ matrices $\{\mathbf{M}_{m}\}_{m=1}^{M}$ as its diagonal blocks. Despite being block-diagonal, matrix $\mathbf{M}$ may still occupy up to $N$ colors. Critically, we do not necessarily have to work with $\mathbf{M}$. Instead, we can permute variables to effect a new measurement matrix that is amenable to more efficient measurements. Suppose we permute vector $\mathbf{v}$ by a permutation matrix $\mathbf{P}$ to get $\mathbf{P}\mathbf{v}$. This entails reordering the nodes of the power system, which is a trivial task that can be performed before solving the OPF. The observable in Lemma 1 can be expressed as
$\sum_{m=1}^{M}\lambda_{m}\mathbf{v}^{\dagger}\mathbf{M}_{m}\mathbf{v}=\sum_{m=1}^{M}\lambda_{m}\mathbf{v}^{\dagger}\mathbf{P}^{\top}(\mathbf{P}\mathbf{M}_{m}\mathbf{P}^{\top})\mathbf{P}\mathbf{v}.$ (27)
@@
Ideally, we would like to design a single $\mathbf{P}$ to minimize the maximum bandwidth across all $\mathbf{P}\mathbf{M}_{m}\mathbf{P}^{\top}$. To this end, we suggest designing $\mathbf{P}$ to minimize the bandwidth of $\mathbf{P}\mathbf{Y}\mathbf{P}^{\top}$, where $\mathbf{Y}$ is the node admittance matrix of the power system. This is because the sparsity pattern of $\mathbf{Y}$ captures the union of sparsity patterns of all $\mathbf{M}_{m}$. Therefore, minimizing the bandwidth of $\mathbf{P}\mathbf{Y}\mathbf{P}^{\top}$ is equivalent to minimizing the maximum bandwidth across all $\mathbf{P}\mathbf{M}_{m}\mathbf{P}^{\top}$.
@@
Unfortunately, minimizing the bandwidth of a matrix is an NP-hard problem even if the matrix is sparse [5]. Nonetheless, heuristic approaches, such as the reverse Cuthill-McKee (RCM) algorithm [6], provide reasonably good solutions in linear complexity $\mathcal{O}(L_{e})=\mathcal{O}(N)$, where $L_{e}$ is the number of edges in the power system graph. The RCM algorithm has been used to reduce the bandwidth of the node admittance matrix $\mathbf{Y}$ in classical computing [52, 53, 54]. Albeit permuting $\mathbf{M}_{m}$’s using the RCM algorithm takes $\mathcal{O}(N)$ operations, the crucial point is that this permutation can be performed once. This is because the power system topology changes infrequently, and such changes only affect a few lines. As long as the network topology remains unchanged, an instance of the OPF is characterized by parameters $b_{m}$, which vary frequently. In contrast to the permutation step, OPF instances over the same topology must be solved repeatedly, every few minutes during real-time operation and potentially millions of times in long-term planning.
@@
Figure 3: The bandwidth of the original and permuted OPF measurement matrices as a function of the logarithmic network size $(\log N)$ for the power system graphs in pglib. Nodes were permuted by the RCM algorithm. Permuted matrices feature patently smaller bandwidths. Upon data fitting and 5-fold cross-validation over different polynomial and exponential functions, the reduced bandwidth was numerically found to scale as $(\log N)^{3}$. The run time of the RCM algorithm (implemented by function scipy.sparse.csgraph.reverse_cuthill_mckee in Python 3.11.9) on the largest power system graph of $N=78,784$ is a few seconds using a MacBook laptop equipped with an M3 Pro processor and 36 GB of RAM.
@@
Figure 4: The number of colors in the original and permuted OPF measurement matrices as a function of the logarithmic network size $(\log N)$ for the power system graphs in pglib. Permuting nodes using the RCM algorithm was numerically found to significantly reduce the number of colors in use. Data fitting and 5-fold cross validation showed that the number of colors scales as $(\log N)^{3}$.
@@
We numerically validated the effect of node permutation on the bandwidth of $\mathbf{M}$. We ran the RCM algorithm on the 66 benchmark power system models of the pglib dataset [55]. Figure 3 shows the bandwidth of $\mathbf{M}$ before and after node permutation. Because the RCM solution depends on the initial node, we conducted 200 Monte Carlo runs per power system, with the initial node drawn uniformly at random per run. Out of the 200 runs, we retained the permutation that yielded the smallest bandwidth. Evidently, permuting nodes reduces the bandwidth of $\mathbf{M}$ dramatically. Upon data fitting and cross-validation, the reduced bandwidth was numerically found to scale as $(\log N)^{3}$. We also empirically tested the effect of node permutation on the number of colors in the OPF measurement matrix. Figure 4 shows that node permutation significantly reduces the number of colors, too. If the bandwidth of $\mathbf{P}\mathbf{M}\mathbf{P}^{\top}$ is $k=\mathcal{O}((\log N)^{3})$ per data fitting in Figure 3, the analysis of [4] upper bounds the number of its colors as $C=\mathcal{O}(k\log N)=\mathcal{O}((\log N)^{4})$. However, the number of colors in $\mathbf{P}\mathbf{M}\mathbf{P}^{\top}$ was found to scale as $(\log N)^{3}$.
@@
We have so far focused on measuring the second observable of the Lagrangian function in (9). The third observable is diagonal and is thus straightforward to measure. The first observable $F_{0}(\boldsymbol{\theta})$, however, is defined over the non-diagonal matrix $\mathbf{M}_{0}$. In general, among all $\mathbf{M}_{m}$’s, matrix $\mathbf{M}_{0}$ could have the most nonzero off-diagonal entries. These entries correspond to power injections from nodes hosting generators and can be $\mathcal{O}(N)$ of those entries. Fortunately, the sparsity pattern of $\mathbf{M}_{0}$ is a subset of the sparsity pattern of $\mathbf{Y}$. Therefore, if nodes are permuted so that $\mathbf{M}$ occupies a few colors $C$, the same permutation works well also for $\mathbf{M}_{0}$. In other words, matrix $\mathbf{P}\mathbf{M}_{0}\mathbf{P}^{\top}$ would occupy $C$ colors or fewer.
@@
Since node permutation yields OPF measurement matrices with much fewer colors, we suggest the following workflow: i) Given a power system model, feed the sparsity pattern of $\mathbf{Y}$ into the RCM algorithm to find a near-optimal node permutation matrix $\mathbf{P}$; ii) Formulate the OPF based on the permuted nodes; iii) Solve the OPF using the VQA of Section 4; and iv) If needed, apply the reverse permutation to the OPF primal solution to recover the original node ordering. It is henceforth assumed that primal variables and measurement matrices have been permuted already. Therefore, we will henceforth use the lighter notation $\mathbf{M}$ and $\mathbf{M}_{m}$ instead of $\mathbf{P}\mathbf{M}\mathbf{P}^{\top}$ and $\mathbf{P}\mathbf{M}_{m}\mathbf{P}^{\top}$.
6.3 Implementation details
@@
Altogether, each PD iteration requires running $\left(2C-1\right)\left(2P+1\right)$ rotated primal circuits and $2Q+1$ dual circuits. Each circuit is run for $S$ samples. According to [13], the number of PQC parameters $(P,Q)$ is expected to be $\tilde{\mathcal{O}}(\log N)$. If the number of colors grows approximately as $C=\mathcal{O}((\log N)^{3})$, each PD iteration in (13) incurs a computational complexity of $\tilde{\mathcal{O}}(\log N)$, which is more efficient compared to $\mathcal{O}(N)$ of the PD iteration in (3) running on a classical computer.
7 Convergence analysis
@@
Proposition 1 shows that the total number of measurement samples scales polynomially with the number of colors $C$, the number of PQC parameters $(P,Q)$, the Lipschitz constant $L$, and the variance bound $\sigma^{2}$.
@@
As discussed in Section 6.3, the number of colors $C$ and the number of PQC parameters $(P,Q)$ grow polynomially in $\log N$. Nonetheless, the dependence of $L$ in (3) and $\sigma^{2}$ in (4) on $\bar{\alpha}$ and $\bar{\beta}$ is unfavorable. This is because the magnitudes of the entries in $\mathbf{v}$ in (QCQP) are constrained to lie in the range of $[0.9,1.1]$, so that $\bar{\alpha}=1.1\sqrt{N}$. Furthermore, as OPF entails equality constraints for nodes connected to loads, each of these equality constraints is converted to two inequality constraints upon (QCQP). Since these inequalities should be active at optimality, the complementary slackness of the Karush–Kuhn–Tucker conditions written for (QCQP) implies that the associated dual variables are generally nonzero [3, Ch. 5]. Because the number of load nodes in real-world power systems is typically proportional to $N$, it follows that $\bar{\beta}$ may scale with $\sqrt{N}$ too. Overall, in practice, parameters $L$ in (3) and $\sigma^{2}$ in (4) could grow as $N$ and $N^{2}$, respectively, which indicates the total number of measurement samples in (40) could scale with $N^{4}$. This large number of measurement samples ultimately presents a challenge to attaining quantum advantage. A similar undesirable dependence of the measurement sample complexity on the scaling prefactors has also been encountered in [57], where the objective function entails expectations multiplied by prefactors exponentially in $\log N$, like the one in (9).
8 Numerical tests
@@
The performance of the proposed doubly variational quantum OPF (Q-OPF) approach was validated using the IEEE 57-node power system benchmark [55]. For this benchmark, we have $N=57$ primal variables and $M=422$ dual variables. Therefore, the number of qubits needed to capture $\mathbf{v}(\boldsymbol{\theta})$ and $\boldsymbol{\lambda}(\boldsymbol{\phi})$ is 6 and 9, respectively. To simplify the implementation, nodes hosting generators were assumed to have no load, i.e., $p^{d}_{n}=q^{d}_{n}=0$ for all $n\in\mathcal{N}_{g}$. Reactive power demands were set so that load nodes have $q^{d}_{n}=0.33\times p^{d}_{n}$ for all $n\in\mathcal{N}_{l}$. We generated 15 different instances of the OPF by scaling original load demands $\{(p^{d}_{n},q^{d}_{n})\}_{n\in\mathcal{N}_{\ell}}$ by factors drawn uniformly distributed within $[0.90,1.05]$, independently per node and problem instance. Each of the 15 OPF instances was solved to global optimality using MATPOWER, a power system toolbox in MATLAB [58].
@@
Model $\mathbf{x}$ [%] $\boldsymbol{\lambda}$ [%] (a) (b) (c)
@@
We first explored the performance of Q-OPF in finding AC OPF solutions. Rather than evaluating errors on OPF solutions in terms of voltage phasors $\mathbf{v}$, we focused on finding the active power injection and voltage magnitude at all generator nodes. This is because power system operators are primarily interested in generator setpoints. Define the vector of generator setpoints $\mathbf{x}=[\mathbf{p}_{g};\mathbf{v}_{g}]^{\top}$ with $\mathbf{p}_{g}$ and $\mathbf{v}_{g}$ collecting generator active power injections and voltage magnitudes, respectively. Generator active power injections can be computed from (44b) given $\mathbf{v}$ and the corresponding load demand $p_{n}^{d}$. In terms of dual variables, we focus on Lagrange multipliers $\lambda_{m}$ associated with power balance and line limit constraints because these are used in electricity markets to compute the so-called locational marginal prices.
@@
We compared solving (QCQP) and (QCQPθ) in terms of the relative errors $\left\|\mathbf{x}-\mathbf{x}^{*}\right\|_{2}/\left\|\mathbf{x}^{*}\right\|_{2}$ and $\left\|\boldsymbol{\lambda}-\boldsymbol{\lambda}^{*}\right\|_{2}/\left\|\boldsymbol{\lambda}^{*}\right\|_{2}$ of the found generator setpoints $\mathbf{x}$ and dual variables $\boldsymbol{\lambda}$ from their global optimal solutions found using MATPOWER. The entries of the dual vector $\boldsymbol{\lambda}$ correspond to the Lagrangian multipliers associated with 280 power balance and line limit constraints. The non-convex (QCQP) was solved classically using the PD iteration in (3) and the EG iteration in (21) over voltage phasors $\mathbf{v}$ and dual variables $\boldsymbol{\lambda}$. Voltages were initialized at the flat voltage profile as $\mathbf{v}^{0}=\mathbf{1}$, while the entries of $\boldsymbol{\lambda}$ were initialized independently at random based on the standard normal distribution and scaled by $2\left|\mathcal{N}_{\ell}\right|$. The related step sizes were set as $\mu_{v}=\mu_{\lambda}=10^{-3}\times 0.9999^{t}$, and PD iterates were terminated when $\left\|\mathbf{v}^{t}-\mathbf{v}^{t-1}\right\|_{2}\leq 10^{-6}$ and $\left\|\boldsymbol{\lambda}^{t}-\boldsymbol{\lambda}^{t-1}\right\|_{2}\leq 10^{-6}$ were met.
Table 1 shows the relative errors for generator setpoints and dual variables attained by solving (QCQPθ) and (QCQP) via the PD and EG iterations. Among the four models, solving (QCQPθ) using the EG iteration achieved the smallest relative errors on $\mathbf{x}$ and $\boldsymbol{\lambda}$. Comparing algorithms when solving either (QCQPθ) or (QCQP), the EG method consistently outperforms the PD method. Similarly, fixing the algorithm to be either EG or PD, solving (QCQPθ) using a hybrid quantum-classical computer offers smaller errors compared to solving (QCQP) using the PD method.
@@
Feasibility. Does the power system safely operate under the generator setpoints computed by Q-OPFs? To answer this question, for each instance, given the input load demands and the setpoints $\mathbf{x}$ obtained by Q-OPF, we found an AC power flow solution via MATPOWER. The obtained voltages were used to evaluate all 222 inequality constraints in (44). Violations in line current and generator constraints were normalized by their maximum limits. We employed three constraint violation metrics: a) the average number of constraint violations surpassing a normalized magnitude of $10^{-6}$ across OPF instances; b) the maximum constraint violation incurred across instances; c) the violations averaged over all constraints and instances. As shown in Table 1, solving (QCQPθ) using the EG iteration attains the best constraint violation metrics among the tested models.

### v2

@@ V.2 Node permutation for qubit-efficient measurements
To solve the doubly parameterized QCQP/SDP, we need to evaluate the Lagrangian function in (10) by measuring the observable in Lemma 1 so that we can compute its gradients over $(\boldsymbol{\theta},\alpha;\boldsymbol{\phi},\beta)$. We next adopt the XBM protocol to efficiently measure the related observables in (10).
@@
According to (18), observable matrix $\mathbf{M}$ has the $N\times N$ matrices $\{\mathbf{M}_{m}\}_{m=1}^{M}$ as its diagonal blocks. Despite being block-diagonal, matrix $\mathbf{M}$ may still occupy up to $N$ colors. Critically, we do not necessarily have to work with $\mathbf{M}$. Instead, we can permute variables to effect a new measurement matrix that is amenable to more efficient measurements. Suppose we permute vector $\mathbf{x}$ by a permutation matrix $\mathbf{P}$ to get $\mathbf{P}\mathbf{x}$. This entails reordering the nodes of the underlying graph, which is a trivial task that can be performed before solving (QCQP). The observable in Lemma 1 can be expressed as
$\sum_{m=1}^{M}\lambda_{m}\mathbf{x}^{\dagger}\mathbf{M}_{m}\mathbf{x}=\sum_{m=1}^{M}\lambda_{m}\mathbf{x}^{\dagger}\mathbf{P}^{\top}(\mathbf{P}\mathbf{M}_{m}\mathbf{P}^{\top})\mathbf{P}\mathbf{x}.$ (28)
@@
Ideally, we would like to design a single $\mathbf{P}$ to minimize the maximum bandwidth across all $\mathbf{P}\mathbf{M}_{m}\mathbf{P}^{\top}$. For many conic programs defined over sparse graphs, the constraint matrices $\mathbf{M}_{m}$ are dictated by a common underlying graph. In such problems, there often exists a representative matrix $\mathbf{Y}$, for instance, the graph Laplacian matrix, whose sparsity pattern captures the union of the sparsity patterns of all $\mathbf{M}_{m}$. Therefore, we propose designing $\mathbf{P}$ to minimize the bandwidth of $\mathbf{P}\mathbf{Y}\mathbf{P}^{\top}$. This is because minimizing the bandwidth of $\mathbf{P}\mathbf{Y}\mathbf{P}^{\top}$ is equivalent to minimizing the maximum bandwidth across all $\mathbf{P}\mathbf{M}_{m}\mathbf{P}^{\top}$. In the OPF problem (detailed in Section VII), the representative matrix is the node admittance matrix $\mathbf{Y}$ of the power network. We henceforth assume such a representative matrix $\mathbf{Y}$ is available or can be constructed for the problem at hand.
@@
Unfortunately, minimizing the bandwidth of a matrix is an NP-hard problem even if the matrix is sparse [18]. Nonetheless, heuristic approaches, such as the reverse Cuthill-McKee (RCM) algorithm [14], provide reasonably good solutions in linear complexity $\mathcal{O}(L_{e})=\mathcal{O}(N)$, where $L_{e}$ is the number of edges in the underlying graph. The RCM algorithm has been used in classical computing to reduce the bandwidth of sparse matrices arising in various domains, including power systems, where it is applied to the node admittance matrix $\mathbf{Y}$ [55, 20, 13]. Albeit permuting $\mathbf{M}_{m}$’s using the RCM algorithm takes $\mathcal{O}(N)$ operations, the crucial point is that this permutation can be performed once for a given graph topology. In applications where the graph structure remains fixed while other problem parameters vary, the same permutation can be reused across multiple problem instances. For example, in the OPF problem, the power system topology changes infrequently and affects only a few edges of the graph. As long as the network topology remains unchanged, an instance of the OPF is characterized by parameters $b_{m}$, which vary frequently. In contrast to the permutation step, OPF instances over the same topology must be solved repeatedly, every few minutes during real-time operation and potentially millions of times in long-term planning.
@@
Figure 3: The bandwidth of the original and permuted measurement matrices as a function of the logarithmic network size $(\log N)$ for the power system graphs in pglib. Nodes were permuted by the RCM algorithm. Permuted matrices feature patently smaller bandwidths. Upon data fitting and 5-fold cross-validation over different polynomial and exponential functions, the reduced bandwidth was numerically found to scale as $(\log N)^{3}$. The run time of the RCM algorithm (implemented by function scipy.sparse.csgraph.reverse_cuthill_mckee in Python 3.11.9) on the largest power system graph of $N=78,784$ is a few seconds using a MacBook laptop equipped with an M3 Pro processor and 36 GB of RAM.
@@
Figure 4: The number of colors in the original and permuted measurement matrices as a function of the logarithmic network size $(\log N)$ for the power system graphs in pglib. Permuting nodes using the RCM algorithm was numerically found to significantly reduce the number of colors in use. Data fitting and 5-fold cross validation showed that the number of colors scales as $(\log N)^{3}$.
@@
We numerically validated the effect of node permutation on the bandwidth of $\mathbf{M}$. We ran the RCM algorithm on the 66 benchmark power system models of the pglib dataset [2]. Figure 3 shows the bandwidth of $\mathbf{M}$ before and after node permutation. Because the RCM solution depends on the initial node, we conducted 200 Monte Carlo runs per power system, with the initial node drawn uniformly at random per run. Out of the 200 runs, we retained the permutation that yielded the smallest bandwidth. Evidently, permuting nodes reduces the bandwidth of $\mathbf{M}$ dramatically. Upon data fitting and cross-validation, the reduced bandwidth was numerically found to scale as $(\log N)^{3}$. We also empirically tested the effect of node permutation on the number of colors in $\mathbf{M}$ for these power systems. Figure 4 shows that node permutation significantly reduces the number of colors, too. If the bandwidth of $\mathbf{P}\mathbf{M}\mathbf{P}^{\top}$ is $k=\mathcal{O}((\log N)^{3})$ per data fitting in Figure 3, the analysis of [33] upper bounds the number of its colors as $C=\mathcal{O}(k\log N)=\mathcal{O}((\log N)^{4})$. However, the number of colors in $\mathbf{P}\mathbf{M}\mathbf{P}^{\top}$ was found to scale as $(\log N)^{3}$.
@@
We have so far focused on measuring the second observable of the Lagrangian function in (10). The third observable is diagonal and is thus straightforward to measure. The first observable $F_{0}(\boldsymbol{\theta})$, however, is defined over the non-diagonal matrix $\mathbf{M}_{0}$. Fortunately, the sparsity pattern of $\mathbf{M}_{0}$ is a subset of the sparsity pattern of $\mathbf{Y}$. Therefore, if nodes are permuted so that $\mathbf{M}$ occupies a few colors $C$, the same permutation works well also for $\mathbf{M}_{0}$. In other words, matrix $\mathbf{P}\mathbf{M}_{0}\mathbf{P}^{\top}$ would occupy $C$ colors or fewer.
@@
Since permuting the underlying graph can yield measurement matrices with much fewer colors, we suggest the following workflow for solving doubly parameterized QCQPs over sparse graphs: i) Identify the underlying graph and construct a representative matrix $\mathbf{Y}$ whose sparsity pattern is the union of the sparsity patterns of all $\mathbf{M}_{m}$. For the OPF, $\mathbf{Y}$ is the node admittance matrix; ii) Feed the sparsity pattern of $\mathbf{Y}$ into the RCM algorithm to find a near-optimal node permutation matrix $\mathbf{P}$; iii) Formulate (QCQP) based on the permuted nodes; iv) Solve (QCQP) using the VQA of Section III; and v) If needed, apply the reverse permutation to the primal solution of (QCQP) to recover the original node ordering. It is henceforth assumed that primal variables and measurement matrices have been permuted already. Therefore, we will henceforth use the lighter notation $\mathbf{M}$ and $\mathbf{M}_{m}$ instead of $\mathbf{P}\mathbf{M}\mathbf{P}^{\top}$ and $\mathbf{P}\mathbf{M}_{m}\mathbf{P}^{\top}$.
V.3 Implementation details
@@
Altogether, each PD iteration requires running $\left(2C-1\right)\left(2P+1\right)$ rotated primal circuits and $2Q+1$ dual circuits. Each circuit is run for $S$ samples. According to [52], the number of PQC parameters $(P,Q)$ is expected to be $\tilde{\mathcal{O}}(\log N)$. If the number of colors grows approximately as $C=\mathcal{O}((\log N)^{3})$, each PD iteration in (14) incurs a computational complexity of $\tilde{\mathcal{O}}(\log N)$, which is more efficient compared to $\mathcal{O}(sN^{2})$ of the PD iteration in (3) running on a classical computer.
VI Convergence analysis
@@
Proposition 1 establishes that the total number of measurement samples scales polynomially with the number of colors $C$, the number of PQC parameters $(P,Q)$, the Lipschitz constant $L$, and the variance bound $\sigma^{2}$. Numerical tests with the OPF in Section V.2 suggest that the number of colors scales as $C=\mathcal{O}((\log N)^{3})$, a scaling we expect to extend to other problems with similar graph sparsity. Additionally, the parameter counts $P$ and $Q$ typically scale as $\tilde{\mathcal{O}}(\log N)$ [52]. The constants $L$ and $\sigma^{2}$ depend on problem-specific bounds $\bar{\alpha}$ and $\bar{\beta}$. If these bounds scale as $\tilde{\mathcal{O}}(\log N)$, then the total number of measurement samples, and hence the overall complexity, becomes $\tilde{\mathcal{O}}(\log N)$, offering a potential advantage over the PD method running on a classical computer. The latter requires $\mathcal{O}(sN^{2})$ operations multiplying with the number of PD iterations in (3). The following section models the OPF as a QCQP and expounds how the sparsity structure of the power network aligns with the measurement protocol presented in Section V.
VII Optimal power flow as a QCQP
$\underaccent{\bar}{p}_{n}\leq p_{n}\leq\bar{p}_{n}\quad\text{and}\quad\underaccent{\bar}{q}_{n}\leq q_{n}\leq\bar{q}_{n}\quad\text{for all}\penalty 10000\ n\in\mathcal{N}.$
$\underaccent{\bar}{v}_{n}^{2}\leq\left|v_{n}\right|^{2}\leq\bar{v}_{n}^{2}\quad\text{for all}\penalty 10000\ n\in\mathcal{N}.$
Line current magnitudes are upper-limited by line capacities as
$\left|i_{\ell}\right|^{2}\leq\bar{i}_{\ell}^{2}\quad\text{for all}\penalty 10000\ \ell\in\mathcal{E}.$
Since all quantities mentioned above are quadratic in $\mathbf{v}$, the OPF can be posed as a quadratically constrained quadratic program (QCQP) over $\mathbf{v}$ as [38, 32]: where $b_{m}$, for all $m$, is a given real-valued parameter determined by lower/upper limits and each $\mathbf{M}_{m}$ is a known Hermitian matrix. For example, if constraint $\underaccent{\bar}{p}_{n}\leq p_{n}$ is indexed as the $m$-th constraint, then $b_{m}=-\underaccent{\bar}{p}_{n}$. The OPF involves quadratic equality constraints, e.g., when $\underaccent{\bar}{p}_{n}=\bar{p}_{n}$ for a subset of nodes $n\in\mathcal{N}$. Each of these equality constraints corresponds to two inequality constraints in (QCQP). Overall, the number of OPF constraints can be several times the network size, such as $M\simeq 8N$.
The precise form of $\mathbf{M}_{0}$ and $\{(\mathbf{M}_{m},b_{m})\}_{m=1}^{M}$ is delineated in Appendix A.1. It is worth stressing that every matrix $\mathbf{M}_{m}$ is highly sparse. This is because every node of a power system is directly connected to only a few other nodes. The number of nonzero entries per matrix $\mathbf{M}_{m}$ is upper bounded by a small constant much smaller than $N$. The $(i,j)$-th off-diagonal entry across every matrix $\mathbf{M}_{m}$ is nonzero only if nodes $i$ and $j$ are connected through a transmission line, that is, if $(i,j)\in\mathcal{E}$. Consequently, the sparsity pattern (location of nonzero entries) for each $\mathbf{M}_{m}$ is determined by a small subset of $\mathcal{E}$. Therefore, the union of the sparsity patterns across all matrices $\mathbf{M}_{m}$ coincides with the sparsity pattern of the nodal admittance matrix $\mathbf{Y}$. The sparsity of the matrices $\mathbf{M}_{m}$} combined with node permutation allows the number of colors $C$ to scale favorably as $\tilde{\mathcal{O}}((\log N)^{3})$, as demonstrated in Section V.3. This efficient coloring enables the measurement protocol of Section V.
@@ VII Optimal power flow as a QCQP
Nonetheless, the overall measurement complexity derived in Proposition 1 also depends on the Lipschitz constant $L$ and the variance bound $\sigma^{2}$ whose scaling is unfavorable for OPF. This is because the magnitudes of the entries in $\mathbf{v}$ in (QCQP) are constrained to lie in the range of $[0.9,1.1]$, so that $\bar{\alpha}=1.1\sqrt{N}$. Furthermore, as OPF entails equality constraints for nodes connected to loads, each of these equality constraints is converted to two inequality constraints upon (QCQP). Since these inequalities should be active at optimality, the complementary slackness of the Karush–Kuhn–Tucker conditions written for (QCQP) implies that the associated dual variables are generally nonzero [8, Ch. 5]. Because the number of load nodes in real-world power systems is typically proportional to $N$, it follows that $\bar{\beta}$ may scale with $\sqrt{N}$ too. Overall, in practice, parameters $L$ in (3) and $\sigma^{2}$ in (4) could grow as $N$ and $N^{2}$, respectively, which indicates the total number of measurement samples in (41) could scale with $N^{4}$. This large number of measurement samples ultimately presents a challenge to attaining quantum advantage. A similar undesirable dependence of the measurement sample complexity on the scaling prefactors has also been encountered in [56], where the objective function entails expectations multiplied by prefactors exponentially in $\log N$, like the one in (10).
VIII Numerical Tests
@@
The performance of the proposed doubly variational quantum QCQP was validated for the OPF problem (hereafter referred to as Q-OPF) using the IEEE 57-node power system benchmark [2]. For this benchmark, we have $N=57$ primal variables and $M=422$ dual variables. Therefore, the number of qubits needed to capture $\mathbf{v}(\boldsymbol{\theta})$ and $\boldsymbol{\lambda}(\boldsymbol{\phi})$ is 6 and 9, respectively. To simplify the implementation, nodes hosting generators were assumed to have no load, i.e., $p^{d}_{n}=q^{d}_{n}=0$ for all $n\in\mathcal{N}_{g}$. Reactive power demands were set so that load nodes have $q^{d}_{n}=0.33\times p^{d}_{n}$ for all $n\in\mathcal{N}_{l}$. We generated 15 different instances of the OPF by scaling original load demands $\{(p^{d}_{n},q^{d}_{n})\}_{n\in\mathcal{N}_{\ell}}$ by factors drawn uniformly distributed within $[0.90,1.05]$, independently per node and problem instance. Each of the 15 OPF instances was solved to global optimality using MATPOWER, a power system toolbox in MATLAB [65].
@@
Model $\mathbf{x}_{g}$ [%] $\boldsymbol{\lambda}$ [%] (a) (b) (c)
@@
We first explored the performance of Q-OPF in finding AC OPF solutions. Rather than evaluating errors on OPF solutions in terms of voltage phasors $\mathbf{v}$, we focused on finding the active power injection and voltage magnitude at all generator nodes. This is because power system operators are primarily interested in generator setpoints. Define the vector of generator setpoints $\mathbf{x}_{g}=[\mathbf{p}_{g};\mathbf{v}_{g}]^{\top}$ with $\mathbf{p}_{g}$ and $\mathbf{v}_{g}$ collecting generator active power injections and voltage magnitudes, respectively. Generator active power injections can be computed from (45b) given $\mathbf{v}$ and the corresponding load demand $p_{n}^{d}$. In terms of dual variables, we focus on Lagrange multipliers $\lambda_{m}$ associated with power balance and line limit constraints because these are used in electricity markets to compute the so-called locational marginal prices.
@@
We compared solving (QCQP) and (QCQPθ) in terms of the relative errors $\left\|\mathbf{x}_{g}-\mathbf{x}_{g}^{*}\right\|_{2}/\left\|\mathbf{x}_{g}^{*}\right\|_{2}$ and $\left\|\boldsymbol{\lambda}-\boldsymbol{\lambda}^{*}\right\|_{2}/\left\|\boldsymbol{\lambda}^{*}\right\|_{2}$ of the found generator setpoints $\mathbf{x}$ and dual variables $\boldsymbol{\lambda}$ from their global optimal solutions found using MATPOWER. The entries of the dual vector $\boldsymbol{\lambda}$ correspond to the Lagrangian multipliers associated with 280 power balance and line limit constraints. The non-convex (QCQP) was solved classically using the PD iteration in (3) and the EG iteration in (22) over voltage phasors $\mathbf{v}$ and dual variables $\boldsymbol{\lambda}$. Voltages were initialized at the flat voltage profile as $\mathbf{v}^{0}=\mathbf{1}$, while the entries of $\boldsymbol{\lambda}$ were initialized independently at random based on the standard normal distribution and scaled by $2\left|\mathcal{N}_{\ell}\right|$. The related step sizes were set as $\mu_{v}=\mu_{\lambda}=10^{-3}\times 0.9999^{t}$, and PD iterates were terminated when $\left\|\mathbf{v}^{t}-\mathbf{v}^{t-1}\right\|_{2}\leq 10^{-6}$ and $\left\|\boldsymbol{\lambda}^{t}-\boldsymbol{\lambda}^{t-1}\right\|_{2}\leq 10^{-6}$ were met.
Table 1 shows the relative errors for generator setpoints and dual variables attained by solving (QCQPθ) and (QCQP) via the PD and EG iterations. Among the four models, solving (QCQPθ) using the EG iteration achieved the smallest relative errors on $\mathbf{x}_{g}$ and $\boldsymbol{\lambda}$. Comparing algorithms when solving either (QCQPθ) or (QCQP), the EG method consistently outperforms the PD method. Similarly, fixing the algorithm to be either EG or PD, solving (QCQPθ) using a hybrid quantum-classical computer offers smaller errors compared to solving (QCQP) using the PD method.
@@
Feasibility. Does the power system safely operate under the generator setpoints computed by Q-OPFs? To answer this question, for each instance, given the input load demands and the setpoints $\mathbf{x}_{g}$ obtained by Q-OPF, we found an AC power flow solution via MATPOWER. The obtained voltages were used to evaluate all 222 inequality constraints in (45). Violations in line current and generator constraints were normalized by their maximum limits. We employed three constraint violation metrics: a) the average number of constraint violations surpassing a normalized magnitude of $10^{-6}$ across OPF instances; b) the maximum constraint violation incurred across instances; c) the violations averaged over all constraints and instances. As shown in Table 1, solving (QCQPθ) using the EG iteration attains the best constraint violation metrics among the tested models.